OpenStudy (anonymous):

f(x)=2^x find the derivative

OpenStudy (lgbasallote):

here's an example that maybe you can follow suppose you are given y = 4^x first take the ln of both sides \[\implies \ln y = x\ln 4\] now you take the derivative of both sides \[\implies \frac 1y y^\prime = \ln 4\] cross multiply \[\implies y^\prime = y\ln 4\] subsittue y \[\implies y^\prime = 4^x \ln 4\] does that help?

OpenStudy (lgbasallote):

or if you want to memorize formulas... \[\huge \frac{\text d}{\text {dx}} (a^\text x) = a^\text x \ln a\]

OpenStudy (lgbasallote):

where a is a constant

OpenStudy (anonymous):

Or.. \[\large log_e{y} = xlog_e4 => y = e^{xln4}\] \[=> \large\frac{dy}{dx} = ln4*e^{xln4} => ln4*4^{x}\]